### v1

@@ Appendix B More Results in Internal Evaluation
Lung 0.660 [0.644, 0.675] 0.727 [0.712, 0.740] 0.811 [0.798, 0.825] 0.399 [0.385, 0.412] 0.373 [0.359, 0.387] 0.068 [0.060, 0.077] 0.813 [0.801, 0.824] $\leq 0.01$
Prostate 0.603 [0.598, 0.609] 0.720 [0.716, 0.724] 0.651 [0.646, 0.657] 0.781 [0.778, 0.784] 0.613 [0.610, 0.616] 0.246 [0.243, 0.249] 0.805 [0.801, 0.809] $\leq 0.01$
Colon 0.545 [0.539, 0.552] 0.438 [0.430, 0.446] 0.660 [0.653, 0.666] 0.443 [0.435, 0.451] 0.371 [0.364, 0.378] 0.407 [0.400, 0.414] 0.727 [0.722, 0.733] $\leq 0.01$
Ovarian 0.607 [0.514, 0.692] 0.417 [0.325, 0.508] 0.634 [0.542, 0.720] 0.493 [0.426, 0.567] 0.446 [0.384, 0.516] 0.713 [0.640, 0.780] 0.684 [0.598, 0.758] -
Esophagus 0.603 [0.537, 0.663] 0.506 [0.441, 0.568] 0.628 [0.564, 0.686] 0.370 [0.317, 0.425] 0.329 [0.279, 0.384] 0.668 [0.611, 0.722] 0.681 [0.623, 0.739] $\leq 0.01$
Bladder 0.613 [0.521, 0.708] 0.472 [0.381, 0.573] 0.629 [0.540, 0.721] 0.365 [0.288, 0.442] 0.334 [0.252, 0.415] 0.700 [0.629, 0.770] 0.675 [0.593, 0.756] -
Testis 0.578 [0.518, 0.639] 0.252 [0.189, 0.316] 0.615 [0.558, 0.672] 0.314 [0.255, 0.372] 0.273 [0.216, 0.332] 0.649 [0.597, 0.704] 0.674 [0.620, 0.733] $\leq 0.01$
Liver 0.555 [0.502, 0.605] 0.465 [0.406, 0.525] 0.628 [0.575, 0.684] 0.389 [0.340, 0.444] 0.326 [0.278, 0.381] 0.664 [0.618, 0.707] 0.658 [0.600, 0.714] -
Pancreatic 0.547 [0.488, 0.598] 0.319 [0.261, 0.377] 0.530 [0.469, 0.585] 0.366 [0.311, 0.417] 0.302 [0.253, 0.351] 0.625 [0.569, 0.672] 0.644 [0.585, 0.695] $\leq 0.01$
Bile duct 0.536 [0.493, 0.576] 0.366 [0.324, 0.410] 0.576 [0.535, 0.615] 0.294 [0.263, 0.328] 0.261 [0.229, 0.294] 0.633 [0.599, 0.669] 0.641 [0.604, 0.678] $\leq 0.01$
Kidney 0.517 [0.442, 0.588] 0.330 [0.262, 0.402] 0.399 [0.326, 0.481] 0.293 [0.238, 0.353] 0.262 [0.204, 0.323] 0.610 [0.541, 0.673] 0.624 [0.554, 0.693] $\leq 0.01$
Adrenal 0.526 [0.473, 0.576] 0.403 [0.354, 0.453] 0.563 [0.512, 0.613] 0.365 [0.319, 0.412] 0.333 [0.288, 0.381] 0.603 [0.553, 0.650] 0.620 [0.567, 0.669] $\leq 0.01$
Cervix 0.560 [0.500, 0.621] 0.322 [0.260, 0.387] 0.567 [0.512, 0.623] 0.313 [0.269, 0.358] 0.259 [0.214, 0.306] 0.608 [0.553, 0.662] 0.617 [0.555, 0.681] $\leq 0.01$
Uterus 0.451 [0.369, 0.526] 0.348 [0.270, 0.426] 0.476 [0.398, 0.555] 0.422 [0.366, 0.481] 0.377 [0.325, 0.432] 0.562 [0.491, 0.632] 0.592 [0.513, 0.663] $\leq 0.01$
Skin 0.443 [0.370, 0.513] 0.290 [0.220, 0.360] 0.491 [0.416, 0.563] 0.332 [0.278, 0.384] 0.288 [0.235, 0.342] 0.579 [0.514, 0.640] 0.589 [0.518, 0.654] $\leq 0.01$
Head neck 0.488 [0.433, 0.543] 0.340 [0.289, 0.394] 0.512 [0.460, 0.561] 0.370 [0.320, 0.420] 0.339 [0.292, 0.386] 0.546 [0.497, 0.593] 0.571 [0.514, 0.625] $\leq 0.01$
Breast 0.400 [0.389, 0.410] 0.362 [0.351, 0.373] 0.449 [0.438, 0.459] 0.456 [0.446, 0.465] 0.409 [0.400, 0.418] 0.288 [0.278, 0.298] 0.564 [0.554, 0.574] $\leq 0.01$
Thyroid 0.460 [0.396, 0.524] 0.232 [0.184, 0.289] 0.521 [0.463, 0.579] 0.364 [0.307, 0.418] 0.325 [0.270, 0.377] 0.576 [0.522, 0.630] 0.556 [0.494, 0.615] -
Stomach 0.434 [0.366, 0.508] 0.320 [0.253, 0.393] 0.482 [0.410, 0.552] 0.380 [0.318, 0.446] 0.315 [0.257, 0.377] 0.530 [0.459, 0.598] 0.528 [0.448, 0.599] -
Multiple 0.168 [0.163, 0.173] 0.112 [0.108, 0.116] 0.135 [0.130, 0.139] 0.221 [0.216, 0.226] 0.204 [0.200, 0.208] 0.084 [0.081, 0.086] 0.444 [0.438, 0.449] $\leq 0.01$
@@
Tissue 0.547 [0.543, 0.551] 0.609 [0.604, 0.613] 0.570 [0.566, 0.575] 0.678 [0.675, 0.682] 0.551 [0.549, 0.555] 0.240 [0.237, 0.242] 0.747 [0.744, 0.751] $\leq 0.01$
Nuclei 0.452 [0.445, 0.458] 0.322 [0.316, 0.329] 0.547 [0.541, 0.554] 0.284 [0.278, 0.289] 0.252 [0.247, 0.257] 0.395 [0.388, 0.402] 0.614 [0.607, 0.620] $\leq 0.01$
Cell 0.067 [0.063, 0.071] 0.016 [0.014, 0.019] 0.112 [0.106, 0.117] 0.121 [0.116, 0.127] 0.100 [0.096, 0.105] 0.048 [0.045, 0.050] 0.382 [0.375, 0.388] $\leq 0.01$
@@
Table 15: Dice scores for segmentation results of each category on the internal datasets of PathSeg. Bold indicates the best result and underline indicates the second best. 95% confidence interval (CI) is included in square brackets.
@@
Figure 12: Prompts of Large Language Models (LLMs) that are used to generate lists of text inputs for PathSegmentor.
Appendix C More Results in External Evaluation
@@
External Category Training Dataset Internal Category DeepLabV3+ SAM-Path nnU-Net
@@
Colon-Cell-Malignant epithelial WSSS4LUAD Tumor epithelial tissue 0.002 [0.001, 0.003] 0.004 [0.001, 0.006] 0.003 [0.000, 0.005]
Colon-Cell-Endothelial SegPath Endothelial 0.000 [0.000, 0.000] 0.000 [0.000, 0.000] 0.000 [0.000, 0.000]
TNBC Endothelial 0.005 [-0.002, 0.010] 0.000 [0.000, 0.000] 0.006 [-0.004, 0.013]
Colon-Cell-Fibroblast NuCLS Fibroblast 0.011 [0.002, 0.018] 0.033 [0.015, 0.049] 0.141 [0.093, 0.187]
TNBC Fibroblast 0.014 [0.007, 0.020] 0.000 [0.000, 0.000] 0.093 [0.054, 0.128]
Colon-Cell-Healthy epithelial MoNuSAC2020 Epithelial 0.002 [-0.002, 0.003] 0.109 [-0.041, 0.215] 0.159 [0.045, 0.272]
PanNuke Epithelial 0.000 [0.000, 0.000] 0.000 [0.000, 0.000] 0.000 [0.000, 0.000]
SegPath Epithelial 0.000 [0.000, 0.000] 0.270 [0.224, 0.318] 0.087 [-0.003, 0.168]
TNBC Epithelial 0.000 [0.000, 0.000] 0.000 [0.000, 0.000] 0.004 [-0.003, 0.008]
Colon-Cell-Inflammatory PanNuke Inflammatory 0.000 [0.000, 0.000] 0.000 [0.000, 0.000] 0.011 [0.001, 0.018]
Colon-Cell-Muscle SegPath Smooth muscle cells 0.000 [0.000, 0.000] 0.089 [0.056, 0.121] 0.000 [0.000, 0.000]
@@
Table 17: Dice scores for segmentation results of each category on the external dataset of PathSeg. Overall results are reported by averaging the Dice scores of all samples. Bold indicates the best result and underline indicates the second best. 95% confidence interval (CI) is included in square brackets. P-values are computed using one-sided, paired Student’s $t$-tests between PathSegmentor and the second-best models.
Category Specialized Model Segmentation Foundation Model P-value
Malignant epithelial 0.002 [0.001, 0.003] 0.004 [0.002, 0.007] 0.003 [0.001, 0.006] 0.462 [0.409, 0.524] 0.406 [0.366, 0.452] 0.615 [0.528, 0.689] 0.624 [0.526, 0.702] $\leq 0.01$
Healthy epithelial 0.002 [0.000, 0.005] 0.270 [0.217, 0.316] 0.159 [0.046, 0.272] 0.265 [0.218, 0.335] 0.221 [0.185, 0.252] 0.383 [0.297, 0.463] 0.596 [0.452, 0.715] $\leq 0.01$
Inflammatory 0.000 [0.000, 0.000] 0.000 [0.000, 0.000] 0.011 [0.002, 0.025] 0.127 [0.067, 0.204] 0.107 [0.060, 0.166] 0.197 [0.131, 0.265] 0.301 [0.217, 0.382] $\leq 0.01$
Muscle 0.000 [0.000, 0.000] 0.089 [0.060, 0.119] 0.000 [0.000, 0.000] 0.111 [0.093, 0.130] 0.114 [0.094, 0.131] 0.267 [0.113, 0.425] 0.301 [0.178, 0.432] $\leq 0.01$
Fibroblast 0.014 [0.008, 0.021] 0.033 [0.017, 0.052] 0.141 [0.095, 0.183] 0.093 [0.068, 0.120] 0.090 [0.066, 0.116] 0.065 [0.036, 0.098] 0.148 [0.106, 0.188] $\leq 0.01$
Endothelial 0.005 [0.000, 0.012] 0.000 [0.000, 0.000] 0.006 [0.000, 0.015] 0.121 [0.035, 0.226] 0.070 [0.030, 0.111] 0.032 [0.000, 0.090] 0.034 [0.004, 0.085] -
Overall 0.005 [0.003, 0.008] 0.030 [0.019, 0.044] 0.054 [0.037, 0.075] 0.170 [0.136, 0.206] 0.148 [0.120, 0.177] 0.217 [0.170, 0.263] 0.291 [0.244, 0.335] $\leq 0.01$
@@ Appendix D Explainable Cancer Diagnosis
Data explainability 21 breast-nuclei-epithelial, breast-tissue-tumor, breast-cell-glial, breast-tissue-dcis, breast-nuclei-neoplastic, breast-tissue-necrosis or debris, breast-cell-cancerous, breast-tissue-lymphocytic infiltrate, breast-nuclei-connective, breast-cell-mitosis, breast-tissue-stroma, breast-tissue-fat, breast-nuclei-lymphocyte, breast-nuclei-inflammatory, breast-tissue-plasma cells, breast-nuclei-neutrophil, breast-nuclei-fibroblast, breast-tissue-blood vessel, breast-nuclei-mitotic figure, breast-nuclei-ductal epithelium, breast-nuclei-plasma cells.
Model explainability 12 breast-cell-cancerous, breast-nuclei-ductal epithelium, breast-nuclei-epithelial, breast-cell-mitosis, breast-nuclei-neoplastic, breast-tissue-tumor, breast-tissue-stroma, breast-nuclei-fibroblast, breast-tissue-lymphocytic infiltrate, breast-tissue-necrosis or debris, breast-tissue-dcis, breast-nuclei-mitotic figure.
Visual explainability 22 breast-nuclei-epithelial, breast-tissue-tumor, breast-cell-glial, breast-tissue-dcis, breast-nuclei-neoplastic, breast-tissue-necrosis or debris, breast-cell-cancerous, breast-tissue-lymphocytic infiltrate, breast-nuclei-connective, breast-cell-mitosis, breast-tissue-stroma, breast-tissue-fat, breast-nuclei-lymphocyte, breast-nuclei-inflammatory, breast-tissue-plasma cells, breast-nuclei-neutrophil, breast-nuclei-fibroblast, breast-tissue-blood vessel, breast-nuclei-mitotic figure, breast-nuclei-ductal epithelium, breast-nuclei-plasma cells, other.

### v2

@@ Appendix A Supplementary Material
DeepLabV3+ SAM-Path nnU-Net
MedSAM
(Union Box)
SAM-Med2D
(Union Box)
BiomedParse PathSegmentor
Lung
0.660
[0.644, 0.675]
0.727
[0.712, 0.740]
0.811
[0.798, 0.825]
0.399
[0.385, 0.412]
0.373
[0.359, 0.387]
0.068
[0.060, 0.077]
0.813
[0.801, 0.824]
$\leq 0.01$
Prostate
0.603
[0.598, 0.609]
0.720
[0.716, 0.724]
0.651
[0.646, 0.657]
0.781
[0.778, 0.784]
0.613
[0.610, 0.616]
0.246
[0.243, 0.249]
0.805
[0.801, 0.809]
$\leq 0.01$
Colon
0.545
[0.539, 0.552]
0.438
[0.430, 0.446]
0.660
[0.653, 0.666]
0.443
[0.435, 0.451]
0.371
[0.364, 0.378]
0.407
[0.400, 0.414]
0.727
[0.722, 0.733]
$\leq 0.01$
Ovarian
0.607
[0.514, 0.692]
0.417
[0.325, 0.508]
0.634
[0.542, 0.720]
0.493
[0.426, 0.567]
0.446
[0.384, 0.516]
0.713
[0.640, 0.780]
0.684
[0.598, 0.758]
Esophagus
0.603
[0.537, 0.663]
0.506
[0.441, 0.568]
0.628
[0.564, 0.686]
0.370
[0.317, 0.425]
0.329
[0.279, 0.384]
0.668
[0.611, 0.722]
0.681
[0.623, 0.739]
$\leq 0.01$
Bladder
0.613
[0.521, 0.708]
0.472
[0.381, 0.573]
0.629
[0.540, 0.721]
0.365
[0.288, 0.442]
0.334
[0.252, 0.415]
0.700
[0.629, 0.770]
0.675
[0.593, 0.756]
Testis
0.578
[0.518, 0.639]
0.252
[0.189, 0.316]
0.615
[0.558, 0.672]
0.314
[0.255, 0.372]
0.273
[0.216, 0.332]
0.649
[0.597, 0.704]
0.674
[0.620, 0.733]
$\leq 0.01$
Liver
0.555
[0.502, 0.605]
0.465
[0.406, 0.525]
0.628
[0.575, 0.684]
0.389
[0.340, 0.444]
0.326
[0.278, 0.381]
0.664
[0.618, 0.707]
0.658
[0.600, 0.714]
Pancreatic
0.547
[0.488, 0.598]
0.319
[0.261, 0.377]
0.530
[0.469, 0.585]
0.366
[0.311, 0.417]
0.302
[0.253, 0.351]
0.625
[0.569, 0.672]
0.644
[0.585, 0.695]
$\leq 0.01$
Bile duct
0.536
[0.493, 0.576]
0.366
[0.324, 0.410]
0.576
[0.535, 0.615]
0.294
[0.263, 0.328]
0.261
[0.229, 0.294]
0.633
[0.599, 0.669]
0.641
[0.604, 0.678]
$\leq 0.01$
Kidney
0.517
[0.442, 0.588]
0.330
[0.262, 0.402]
0.399
[0.326, 0.481]
0.293
[0.238, 0.353]
0.262
[0.204, 0.323]
0.610
[0.541, 0.673]
0.624
[0.554, 0.693]
$\leq 0.01$
Adrenal
0.526
[0.473, 0.576]
0.403
[0.354, 0.453]
0.563
[0.512, 0.613]
0.365
[0.319, 0.412]
0.333
[0.288, 0.381]
0.603
[0.553, 0.650]
0.620
[0.567, 0.669]
$\leq 0.01$
Cervix
0.560
[0.500, 0.621]
0.322
[0.260, 0.387]
0.567
[0.512, 0.623]
0.313
[0.269, 0.358]
0.259
[0.214, 0.306]
0.608
[0.553, 0.662]
0.617
[0.555, 0.681]
$\leq 0.01$
Uterus
0.451
[0.369, 0.526]
0.348
[0.270, 0.426]
0.476
[0.398, 0.555]
0.422
[0.366, 0.481]
0.377
[0.325, 0.432]
0.562
[0.491, 0.632]
0.592
[0.513, 0.663]
$\leq 0.01$
Skin
0.443
[0.370, 0.513]
0.290
[0.220, 0.360]
0.491
[0.416, 0.563]
0.332
[0.278, 0.384]
0.288
[0.235, 0.342]
0.579
[0.514, 0.640]
0.589
[0.518, 0.654]
$\leq 0.01$
Head neck
0.488
[0.433, 0.543]
0.340
[0.289, 0.394]
0.512
[0.460, 0.561]
0.370
[0.320, 0.420]
0.339
[0.292, 0.386]
0.546
[0.497, 0.593]
0.571
[0.514, 0.625]
$\leq 0.01$
Breast
0.400
[0.389, 0.410]
0.362
[0.351, 0.373]
0.449
[0.438, 0.459]
0.456
[0.446, 0.465]
0.409
[0.400, 0.418]
0.288
[0.278, 0.298]
0.564
[0.554, 0.574]
$\leq 0.01$
Thyroid
0.460
[0.396, 0.524]
0.232
[0.184, 0.289]
0.521
[0.463, 0.579]
0.364
[0.307, 0.418]
0.325
[0.270, 0.377]
0.576
[0.522, 0.630]
0.556
[0.494, 0.615]
Stomach
0.434
[0.366, 0.508]
0.320
[0.253, 0.393]
0.482
[0.410, 0.552]
0.380
[0.318, 0.446]
0.315
[0.257, 0.377]
0.530
[0.459, 0.598]
0.528
[0.448, 0.599]
Multiple
0.168
[0.163, 0.173]
0.112
[0.108, 0.116]
0.135
[0.130, 0.139]
0.221
[0.216, 0.226]
0.204
[0.200, 0.208]
0.084
[0.081, 0.086]
0.444
[0.438, 0.449]
$\leq 0.01$
@@
DeepLabV3+ SAM-Path nnU-Net
MedSAM
(Union Box)
SAM-Med2D
(Union Box)
BiomedParse PathSegmentor
Tissue
0.547
[0.543, 0.551]
0.609
[0.604, 0.613]
0.570
[0.566, 0.575]
0.678
[0.675, 0.682]
0.551
[0.549, 0.555]
0.240
[0.237, 0.242]
0.747
[0.744, 0.751]
$\leq 0.01$
Nuclei
0.452
[0.445, 0.458]
0.322
[0.316, 0.329]
0.547
[0.541, 0.554]
0.284
[0.278, 0.289]
0.252
[0.247, 0.257]
0.395
[0.388, 0.402]
0.614
[0.607, 0.620]
$\leq 0.01$
Cell
0.067
[0.063, 0.071]
0.016
[0.014, 0.019]
0.112
[0.106, 0.117]
0.121
[0.116, 0.127]
0.100
[0.096, 0.105]
0.048
[0.045, 0.050]
0.382
[0.375, 0.388]
$\leq 0.01$
@@
Table 15: Dice scores for segmentation results of each category on the internal datasets of PathSeg. Overall results are reported by averaging the Dice scores of all samples on the internal datasets. Bold indicates the best result and underline indicates the second best. 95% confidence interval (CI) is included in square brackets.
@@
Figure 12: Prompts of Large Language Model (LLM) that are used to generate lists of text inputs for PathSegmentor.
@@
External Dataset External Category Training Dataset Internal Category DeepLabV3+ SAM-Path nnU-Net
@@
CoNSeP Colon-Cell-Malignant epithelial WSSS4LUAD Tumor epithelial tissue
0.002
[0.001, 0.003]
0.004
[0.001, 0.006]
0.003
[0.000, 0.005]
Colon-Cell-Endothelial SegPath Endothelial
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
TNBC Endothelial
0.005
[-0.002, 0.010]
0.000
[0.000, 0.000]
0.006
[-0.004, 0.013]
Colon-Cell-Fibroblast NuCLS Fibroblast
0.011
[0.002, 0.018]
0.033
[0.015, 0.049]
0.141
[0.093, 0.187]
TNBC Fibroblast
0.014
[0.007, 0.020]
0.000
[0.000, 0.000]
0.093
[0.054, 0.128]
Colon-Cell-Healthy epithelial MoNuSAC2020 Epithelial
0.002
[-0.002, 0.003]
0.109
[-0.041, 0.215]
0.159
[0.045, 0.272]
PanNuke Epithelial
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
SegPath Epithelial
0.000
[0.000, 0.000]
0.270
[0.224, 0.318]
0.087
[-0.003, 0.168]
TNBC Epithelial
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
0.004
[-0.003, 0.008]
Colon-Cell-Inflammatory PanNuke Inflammatory
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
0.011
[0.001, 0.018]
Colon-Cell-Muscle SegPath Smooth muscle cells
0.000
[0.000, 0.000]
0.089
[0.056, 0.121]
0.000
[0.000, 0.000]
CPM17 Unspecified-Nuclei-Nuclei Janowczyk Nuclei
0.052
[0.036, 0.067]
0.274
[0.242, 0.306]
0.111
[0.095, 0.129]
CPM15
0.056
[0.030, 0.086]
0.211
[0.157, 0.273]
0.132
[0.092, 0.170]
Kumar
0.067
[0.045, 0.094]
0.170
[0.133, 0.206]
0.083
[0.059, 0.110]
Lizard
0.010
[0.007, 0.015]
0.044
[0.037, 0.051]
0.017
[0.013, 0.022]
@@
Table 17: Dice scores for segmentation results on the external datasets of PathSeg. For the CoNSeP dataset, we present results for the overall dataset as well as for each category separately. For the other datasets, since they include only one nuclei category, we provide results at the dataset level. Bold indicates the best result and underline indicates the second best. 95% confidence interval (CI) is included in square brackets. P-values are computed using one-sided, paired Student’s $t$-tests between PathSegmentor and the second-best models.
Dataset/Category Specialized Model Segmentation Foundation Model P-value
DeepLabV3+ SAM-Path nnU-Net
MedSAM
(Union Box)
SAM-Med2D
(Union Box)
BiomedParse PathSegmentor
CPM17
0.052
[0.036, 0.067]
0.274
[0.242, 0.306]
0.111
[0.095, 0.129]
0.329
[0.299, 0.355]
0.316
[0.292, 0.342]
0.553
[0.485, 0.619]
0.705
[0.670, 0.738]
$\leq 0.01$
CPM15
0.056
[0.030, 0.086]
0.211
[0.157, 0.273]
0.132
[0.092, 0.170]
0.274
[0.234, 0.308]
0.278
[0.244, 0.312]
0.236
[0.087, 0.396]
0.485
[0.365, 0.602]
$\leq 0.01$
Kumar
0.067
[0.045, 0.094]
0.170
[0.133, 0.206]
0.083
[0.059, 0.110]
0.393
[0.351, 0.435]
0.368
[0.326, 0.408]
0.181
[0.093, 0.269]
0.427
[0.387, 0.465]
$\leq 0.01$
Lizard
0.010
[0.007, 0.015]
0.044
[0.037, 0.051]
0.017
[0.013, 0.022]
0.299
[0.286, 0.310]
0.282
[0.271, 0.293]
0.232
[0.212, 0.254]
0.319
[0.304, 0.333]
$\leq 0.01$
CoNSeP
0.005
[0.003, 0.008]
0.030
[0.019, 0.044]
0.054
[0.037, 0.075]
0.170
[0.136, 0.206]
0.148
[0.120, 0.177]
0.217
[0.170, 0.263]
0.291
[0.244, 0.335]
$\leq 0.01$
Malignant epithelial
0.002
[0.001, 0.003]
0.004
[0.002, 0.007]
0.003
[0.001, 0.006]
0.462
[0.409, 0.524]
0.406
[0.366, 0.452]
0.615
[0.528, 0.689]
0.624
[0.526, 0.702]
$\leq 0.01$
Healthy epithelial
0.002
[0.000, 0.005]
0.270
[0.217, 0.316]
0.159
[0.046, 0.272]
0.265
[0.218, 0.335]
0.221
[0.185, 0.252]
0.383
[0.297, 0.463]
0.596
[0.452, 0.715]
$\leq 0.01$
Inflammatory
0.000
[0.000, 0.000]
0.000
[0.000, 0.000]
0.011
[0.002, 0.025]
0.127
[0.067, 0.204]
0.107
[0.060, 0.166]
0.197
[0.131, 0.265]
0.301
[0.217, 0.382]
$\leq 0.01$
Muscle
0.000
[0.000, 0.000]
0.089
[0.060, 0.119]
0.000
[0.000, 0.000]
0.111
[0.093, 0.130]
0.114
[0.094, 0.131]
0.267
[0.113, 0.425]
0.301
[0.178, 0.432]
$\leq 0.01$
Fibroblast
0.014
[0.008, 0.021]
0.033
[0.017, 0.052]
0.141
[0.095, 0.183]
0.093
[0.068, 0.120]
0.090
[0.066, 0.116]
0.065
[0.036, 0.098]
0.148
[0.106, 0.188]
$\leq 0.01$
Endothelial
0.005
[0.000, 0.012]
0.000
[0.000, 0.000]
0.006
[0.000, 0.015]
0.121
[0.035, 0.226]
0.070
[0.030, 0.111]
0.032
[0.000, 0.090]
0.034
[0.004, 0.085]
@@
Feature Importance Estimation 12 breast-cell-cancerous, breast-nuclei-ductal epithelium, breast-nuclei-epithelial, breast-cell-mitosis, breast-nuclei-neoplastic, breast-tissue-tumor, breast-tissue-stroma, breast-nuclei-fibroblast, breast-tissue-lymphocytic infiltrate, breast-tissue-necrosis or debris, breast-tissue-dcis, breast-nuclei-mitotic figure.
Imaging Biomarker Discovery 22 breast-nuclei-epithelial, breast-tissue-tumor, breast-cell-glial, breast-tissue-dcis, breast-nuclei-neoplastic, breast-tissue-necrosis or debris, breast-cell-cancerous, breast-tissue-lymphocytic infiltrate, breast-nuclei-connective, breast-cell-mitosis, breast-tissue-stroma, breast-tissue-fat, breast-nuclei-lymphocyte, breast-nuclei-inflammatory, breast-tissue-plasma cells, breast-nuclei-neutrophil, breast-nuclei-fibroblast, breast-tissue-blood vessel, breast-nuclei-mitotic figure, breast-nuclei-ductal epithelium, breast-nuclei-plasma cells, other.